17.1 g of C2H6 is burned in oxygen. What mass of oxygen was consumed?(Assumed complete combustion). Show th solution.

1
Expert's answer
2022-05-31T01:50:07-0400

Solution:

The molar mass of C2H6 is 30 g/mol

Therefore,

Moles of C2H6 = (17.1 g C2H6) × (1 mol C2H6 / 30 g C2H6) = 0.57 mol C2H6


Balanced chemical equation:

2C2H6 + 7O2 → 4CO2 + 6H2O

According to stoichiometry:

2 mol of C2H6 react with 7 mol of O2

Thus, 0.57 mol of C2H6 react with:

(0.57 mol C2H6) × (7 mol O2 / 2 mol C2H6) = 1.995 mol O2


The molar mass of O2 is 32 g/mol

Therefore,

Mass of O2 = (1.995 mol O2) × (32 g O2 / 1 mol O2) = 63.84 g O2 = 63.8 g O2

Mass of O2 = 63.8 g


Answer: 63.8 grams of oxygen was consumed